if x + 15.73% = 315, find the value of X

Answer by stanbon(75887)

if x + 15.73% = 315, find the value of X
------------
x + (15.73/100) = 315
---
100x + 15.72 = 31500
100x = 31484.28
x = 314.8428
=======================
Cheers,
Stan H.
============